Find out de-broglie wavelength of proton whose kinetic energy is 10KeV.
1
Expert's answer
2020-09-09T10:21:40-0400

solution

given data

mass of proton (m)"=1.67\times10^{-27}kg"

charge of proton (q)="1.6\times10^{-19}C"

1eV="1.6\times10^{-19}V"

kinetic energy of proton (KE)=10KeV


kinetic energy is given by


"KE=\frac{p^2}{2m}" ........eq.1


de-broglie wavelength is given by


"\lambda=\frac{h}{p}" ........eq.2


by using equation 1 and 2 we get


"\lambda=\frac{h}{\sqrt{2mKE}}"


by putting the value of h ,m and KE


"\lambda=\frac{6.625\times10^{-34}}{\sqrt{2\times1.67\times10^{-27}\times10^4\times1.6\times10^{-19}}}"


"\lambda=\frac{6.625\times10^{-13}}{\sqrt{2}\times1.6}"

"\lambda=2.93\times10^{-13}m"


this is comparable of size relative order to size of nucleus.